Megan bought x pounds of coffee that costs 3 dollars per pound and 18 pounds of coffee at 2.5 per pound. Find the total number of coffee purchased if the average cost per pound of both types together is 2.85
Answer by Boreal(15235)

x*3=3x cost of 1 type. units are $
18*2.50=45 of other type
(18+x)*2.85=51.3+2.85x
3x+45+51.3+2.85x, the amount of the first two equals the total
0.15x=6.3
x=42 pounds
that cost is $126
18 pounds costs $45
60 pounds costs $171
171/60=$2.85
Answer is 60 pounds